OpenStudy (anonymous):

What is the area of a circle with a diameter of 4 inches? Choice 'A' A ≈ 12.52 inches^3 Choice 'B' A = 12.46 inches^2 Choice 'C' A = 12.56 inches^3 Choice 'D' A ≈ 12.56 inches^2

OpenStudy (anonymous):

d

OpenStudy (angela210793):

A=pir^2=pi(d/2)^2=pi*(4/2)^2=4pi